Each answer choice is an answer to one of the Q’s.

media%2F891%2F891b238d-361b-464a-bbf7-19
Answer #1

Q1

Net Revenue

=gross revenue-return and allowance

=536000-32800

=$503200

Q2

CGS=Beginning + purchase -ending Inventory

=$125900+$256800-$125200

=$257500

Q3

Gross Profit =-Net Revenue-CGS=$503200-257500=$245700

Q4

Net Income =Gross profit-Operating exp=247500-128100=$119400

Q5

Current ratio =Current assets/current Liabilities=$114000/69000=1.65

Q6

Acid test ratio=(Current Asset-Inventory)/Current Liabilities =(114000-13900)/69000=1.45

Q7

Average collection Period=360/(Net Sales/Account receivable) =360/(136500/25500)=67.25

Q8

Profit Margin=Net income.net revenue =28000/136500=.2051